OpenStudy (anonymous):

Can Somebody Help? 5(8e^2x - 3)^3=625

OpenStudy (anonymous):

We'll divide by 5 both sides: 5(8e^2x - 3)^3 = 625 (8e^2x - 3)^3 = 125 We'll take cube root both sides: 8e^2x - 3 = 5 We'll add 3 both sides: 8e^2x = 8 We'll divide by 8: e^2x = 1 We'll take natural logarithms both sides: ln e^2x = ln 1 We'll apply the power property of logarithms: 2x ln e = ln 1 But ln e =1 and ln 1 = 0. 2x = 0 The solution of the equation is x = 0.

OpenStudy (anonymous):

basically, this is 8e^2x-3=5, giving, e^2x=1, 2x=Ln(1)=0, x=0

OpenStudy (anonymous):

(8e^2x - 3)^3=125( but125=5^3) so cancelling cube 8e^2x - 3=5 8e^2x =8 e^2x =1 2x=log 1 x=0

jhonyy9 (jhonyy9):

do you understand it ?

jhonyy9 (jhonyy9):

so i think that is right

OpenStudy (anonymous):

Yes I understand

jhonyy9 (jhonyy9):

ok